OpenStudy (anonymous):

In a rectanglular coordinate system, what is the area of a triangle whose vertices whose vertices have the coordinates (4,0), (6, 3) adn (6 , -3)

OpenStudy (anonymous):

area of triangle is = 1/2 * base * height please plot the vertices on a graph it will be easy to calculate

OpenStudy (anonymous):

here height will be 2 units and base will be 6 units so arae will be 6 square units

OpenStudy (anonymous):

hey akankha,This question can be solved by calculating the eucleadian distance between the points x and y so can you care to specify the vertices(Base,hpo,perpendicular)..please rporvide a simple diagram

OpenStudy (anonymous):

here's the way: lets say two vertices of the base be \[V _{B1} \] and \[V _{B2}\] where\[ V_{B1} = [x _{1},y _{1}]\] \[V _{B2} = [x _{2},y _{2}]\]  Now calculate the length of base: \[Base=Distance \between V_{B1} and V _{B2}\] Distance between two points in cartesian/euclidean space : = \[D = \sqrt{(x _{1} - y _{1})^{2} + (x _{2} - y _{2})^{2}}\] by using that formula we can calculate the length of base and height . then you can plug them into : \[Area \Delta = \frac{ 1 }{ 2 } \times base \times height \] And voila,there comes your area of the triangle. NOte:you have to figure it out how to calculate the length of height yourself(I already equipped you with the artillery required(believe me it will be fun))

OpenStudy (anonymous):